OpenStudy (anonymous):

HELP! Angle R and Angle T are complementary. If m angle R = 3x and m angle T = 2x, what is the value of x?

OpenStudy (anonymous):

@Jhannybean @shubhamsrg Help???!!!

OpenStudy (anonymous):

complementary angels means that the sum of those angles must be 90 degrees. hence R + T =90 then replace the given values of R and T in the above equation and solve for x: R +T =90 3x + 2x=90 5x=90 x=90/5=18